OpenStudy (anonymous):

Explain how to solve the equation 3t + 11 = –10. Show that your solution works.

OpenStudy (anonymous):

t = -7 Can't show :( . . . .. ... It's a trade secret.

OpenStudy (anonymous):

Like a sir :P

OpenStudy (tony123):

3t +11=-10 first you subract 11 from both sides to get 3t=-21 then divide both sides by 3 to get t=-7

OpenStudy (anonymous):

Now I'll have to show it... 3t + 11 = -10 3t = -11 -10 3t = -21 t = -21/3 = -7 \[\large\mathsf{\color{red}{\text{LIKE A SIR}}}\]

OpenStudy (anonymous):

Thank you both :)